# Heart Failure Prediction using Modal Decomposition and Masked Autoencoders for Scarce Echocardiography Databases

### v1

@@ 2.2 Heart Failure Prediction
In the second phase, Modal Decomposition-based Data Transform, the same process as in the Modal Decomposition-based Data Generation phase described in Subsection 2.1 is performed. This is, images with less noise and more discriminative features than the original echocardiography data are obtained by sequentially applying the SVD and the HODMD algorithms, increasing the differentiation between the heart states and improving the heart failure prediction performance. Note that, in this case, the resulting images from this phase are exclusively taken as input of the deep neural network for the prediction of time of a heart failure of the corresponding sequence. Therefore, this phase only acts as a feature extractor. However, the Modal Decomposition-based Data Generation phase leverages the output of each applied algorithm to enlarge the training database $D_{train}$, directly fed to the deep neural network for training, containing more discriminative samples for the differentiation between the heart states, and addressing the limited availability of high-quality samples usual in the medicine field. Different types of data resulting from the Modal Decomposition-based Data Transform phase have been tested as input of the deep neural network for the prediction of heart failure times in Section 4. The use of HODMD modes for test tends to give the best heart failure time prediction performances. This demonstrates the efficacy of these techniques to extract more discriminative features associated with the different heart states from echocardiography images and to filter out the high noise inherent, leading to more accurate heart failure time predictions.
@@
The third phase, Deep Neural Network-based Heart Failure Prediction, computes the heart failure time prediction for each image generated in the previous phase, Modal Decomposition-based Data Transform. For this purpose, a deep neural network architecture based on the ViT is used. It has been trained using a self-supervised learning approach which improves the heart failure time prediction performance, even with scarce datasets (Das et al. (2024)). To the best of the authors’ knowledge, self-supervised learning has been barely explored in the related literature about heart failure prediction in echocardiography images, even less Masked Autoencoders (MAE) (He et al. (2022)). Moreover, standard architectures have been proposed as the general trend with no specific adaptations nor approaches to deal with this typical scenario in the medicine field of having a low number of high-quality samples. The training approach used involves two tasks: the Self-supervised Auxiliary Task, which aids the training for the regression task about heart failure time prediction, precisely the one of interest. This allows to increase the locality inductive bias of the ViT, and so to improve the performance on heart failure prediction in the usual scenario of scarse datasets. In addition, this training approach further reduces the dependence on very large training databases with high quality, which are costly to elaborate and which require very hard specialized work. As will be seen in Subsection 4.2, the proposed deep neural network trained with this self-supervised learning approach achieves a better performance than several ViT architectures and Convolutional Neural Networks (CNN), allowing to reduce computational resources.
@@
The architecture is represented in Fig. 5 together with the used training approach based on self-supervised learning. Two tasks are jointly learned at the same time: the self-supervised Auxiliary Task, and the Regression Task. The first one is aimed to reconstruct the missing patches from masked images, improving the discriminating capability of the deep neural network aimed to predict heart failure times, i.e., the one devoted to the Regression Task. Precisely, the weights of the Transformer Encoders used for both tasks are shared, allowing this joint training. Note that any self-supervised learning approach can be used, but the Masked Autoencoder (MAE) (He et al. (2022)) has been employed due to its popularity and superior performance (Das et al. (2024)). This training approach based on joint learning is different to the usual approach of sequentially performing the pretraining for reconstruction of the masked patches, and then the fine-tuning, using the same dataset, for the task of interest (He et al. (2022)). That is, not only improves the discriminating capability and feature learning of the deep neural network, but also allows to speed up the training process by learning both tasks at the same time. The common input for the two tasks is an image representing a heart in a determined state. As a result of the previous phase, Modal Decomposition-based Data Transform, this input image can be a reconstruction, or a mode obtained with the SVD or the HODMD algorithms. This fact makes the deep neural network to have the advantage of being able to directly use images as input, in the form of original echocardiography data (in case of not using the Modal Decomposition-based Data Transform phase), reconstructions, or modes obtained with the modal decomposition techniques. However, as said previously, using the modes obtained with the HODMD algorithm as input for the deep neural network has led to the best performance. The output is a prediction which represents the estimated time in which a heart failure will happen. The ViT, as a deep neural network, is the one which addresses the non-linear relationships which could be present between the input and the output, in this case, between echocardiography images and heart failure times, regardless of whether the SVD and the HODMD algorithms have been employed or not. However, as will be seen in Section 4, the use of the SVD and HODMD algorithms for both feature extraction and data augmentation, and the fact that the proposed ViT is specifically adapted to scarce databases of echocardiography images, allow this proposed deep neural network to effectively model the relationship existing between echocardiography images and heart failure times. In addition, as already seen in Subsubsection 2.1.1, the HODMD algorithm, formulated with linear operators, is an approximation of the Koopman operator, allowing to deal with the non-linearities existing in the echocardiography imagery.
@@
The processing of the input for the Regression Task (the one of interest), for heart failure time prediction closely follows the standard ViT architecture. That is, a division in non-overlapping patches is firstly performed. Next, the spatial dimension of the patches is flattened, followed by a normalization layer and a linear projection. Before processing the patches with the Transformer Encoder, positional information and a regression token, with a similar role to the class token, are added.
@@
Figure 5: Architecture of the proposed deep neural network with the training approach based on self-supervised learning. (a) The input: an image from an echocardiography video sequence, in the form of an original echocardiography sample, a mode, or a reconstruction obtained with the SVD or the HODMD algorithms. (b) The self-supervised Auxiliary Task, aimed to reconstruct the missing patches from the masked image, aiding the training of the ViT for the regression task. (c) The Regression Task, the one of interest, for heart failure time prediction. (d) The predicted heart failure time of the input image.
@@
The backbone which processes the visual tokens is based on a Transformer Encoder, whose weights are shared with the Transformer Encoder used for the Self-supervised Auxiliary Task, learned at the same time. This has been done with twelve stacked Transformer blocks (so $B_{E}=12$). Each Transformer block includes a multi-head attention layer, with three heads, projection dimension $192$, and without dropout.
@@
An empty branch (skipped connection) and a Multi-layer Perceptron (MLP) with two units follows the multi-head attention layer, each one composed of a fully connected layer with the Gaussian Error Linear Unit (GELU) activation function. The use of skip connections addresses the vanishing gradient problem, inherent to deep neural networks, to considerably improve training convergence. The MLP ratio used is $4$, that is, the number of hidden features are four times the amount of input features. The Transformer block finally introduces a skip connection, connecting the output of the previous one to the output of this MLP. After all the $B_{E}$ Transformer blocks, features are normalized and introduced into a regression head, whose output is the estimation of the time in which a heart failure will happen, according to the input image.
@@
Regarding the self-supervised Auxiliary Task, the processing is based on the MAE (He et al. (2022)). That is, first, patches from the input image are randomly masked before the tokenization. After that, the same Transformer Encoder as the one for the Regression Task is used. Precisely, the weights between both Transformer Encoders are shared. Therefore, during training, these weights are adapted to both tasks, improving the discriminating capability of the ViT used for heart failure time prediction, even with scarce databases. After the $B_{E}$ Transformer blocks, a Transformer Decoder is used for the reconstruction of the missing patches, much shallower than the Encoder. This Decoder has been designed by first incorporating a decoder embedding which adapts the projection dimension to $128$ via a fully connected layer. After adding the masked tokens and the position embedding, two stacked Transformer blocks have been used ($B_{D}=2$). Each of these blocks includes a multi-head attention layer with $16$ heads. Similarly, empty branches and an MLP with two units and MLP ratio of 4, each one with the GELU activation function, have been also employed. After all the $B_{D}$ Transformer blocks, features are normalized, the regression token is removed, and a decoder head is used for the reconstruction of the missing patches.
@@
The joint learning on both tasks is expressed with the following final convex loss function:
@@
where $L_{reg}$ is the regression loss between the estimated heart failure time and the real one; $L_{SSAT}$ is the loss component between the original and reconstructed image, computed only for the masked patches (He et al. (2022)), and $\alpha$ is the loss scaling factor, weighting the importance of each task. Both loss functions are based on the Mean Square Error (MSE). Note that, after training, for the Deep Neural Network-based Heart Failure Prediction phase, only the branch concerning the Regression Task (indicated as (c) in Fig. 5) is used.
@@
The deep neural network uses the annotated cardiac database of images, $D_{train}$, obtained in the Cardiac Database Creation stage described in Subsection 2.1 for training. Concerning the optimization technique, the AdamW algorithm has been used. The employed learning rate strategy is based on the warm-up cosine policy (Lee et al. (2021)) that computes the learning rate at iteration $i$ by the expression $\lambda_{i}=0.5\times\lambda_{t}\times(1+cos(\pi\times(i-N_{w})/(N_{iter}-N_{w})))$, where $N_{iter}$ is the maximum number of iterations. The target learning rate is fixed to $\lambda_{t}=2.5\mathrm{e}{-4}$, and the warmup steps $N_{w}$ to $5$. The momentum and weight decay are set to 0.9 and 0.05, respectively. The batch size has been set to 64 images due to the limitations of the physical memory of the GPUs available (Nvidia A100 and Nvidia RTX A4500). About the database division, a splitting scheme of $60\%-20\%$ for training and validation, respectively, has been adopted, balancing the cardiac categories to ensure that every set contains approximately the same number of samples for each heart state. The batch generation scheme includes data augmentation techniques based on geometric transformations (namely: resizing, random horizontal flips, and random erasing), applied to the echocardiography data. In this way, the model becomes more robust to different perspectives acquired of the heart. In addition, these data augmentation techniques complement to the Modal Decomposition-based Data Generation phase already explained in Subsection 2.1, marginally improving the performance compared to the modal decomposition.
@@
In the fourth and last phase, Fusion of Heart Failure Predictions, an estimation of the time of happening a heart failure in the input test sequence is determined by combining multiple predictions. This process is illustrated in Fig. 6. Each image $i$ from the $N+1$ images comprising the video sequence used as input of the previous phase, Deep Neural Network-based Heart Failure Prediction, has associated a time prediction $P_{i}$, that is, the estimation of the age of the patient in which a heart failure will happen. The set of time predictions $\{P_{i}\}$ is then used to compute a unique time prediction $\hat{P}$. For this purpose, the average has been used. As a result, an estimation of the time in which a heart failure will happen is obtained. Note that relying on a sequence of echocardiography images instead of a single image to predict the time of heart failures is more robust, as using more data can help doctors to make more confident heart prognoses. In addition, certain heart pathologies might lead to arrhythmias or patterns in the heart cycles which, therefore, constitute temporal information which can be obtained from sequences, and which could be valuable to accurately predict heart failures.
@@ 3 Database
The heart failure prediction system here proposed has been tested and validated using a database of echocardiography images. It is composed of video sequences representing the following heart states: Control (CTL), Obesity (OB), and Systemic Hypertension (SH). This database has been elaborated in a collaboration with the Centro Nacional de Investigaciones Cardiovasculares (CNIC). A sample image from each of these heart states is shown in Fig. 7. As can be seen, echocardiography imagery inherently has much noise, as expected in this modality of data. In addition, different perspectives are taken, increasing the complexity of the heart failure prediction task. Only the area of the heart, i.e., the region of interest (ROI), is taken for the prediction system, as specified in the Data Homogenization phase described in Subsection 2.1. The obtained heart areas have different resolutions, showing the heterogeneity of the echocardiography imagery: in mean and standard deviation, the spatial resolution is $(675.68\pm 47.28)\times(583.15\pm 0.49)$ pixels. However, the aspect factor among the images barely changes and is close to 1, so these are practically square: $1.16\pm 0.08$. Therefore, resizing to square images for the deep neural networks barely deforms the heart areas, and does not suppose a degradation of the heart failure prediction performance, as preserving the spatial features characterizing the different heart states.
@@
A summary of the main characteristics of the cardiac database is presented in Table 1. It includes information about the times in which heart failures happen. An average and a standard deviation of the age of the heart failure event is represented for each heart state. As can be seen, unhealthy classes (OB, and SH) tend to have earlier heart failures than in the case of healthy hearts (CTL), as expected. This cardiac database is composed of the original echocardiography images, and the associated SVD-based data and the HODMD-based data. These data have been generated as a result of the Modal Decomposition-based Data Generation phase described in Subsection 2.1. That is, the HODMD-based data were obtained after applying the HODMD algorithm on the SVD reconstructions of the original images. Note that the total number of original snapshots equals the number of associated reconstructed ones after applying the SVD algorithm, and also the amount of reconstructions obtained after using the HODMD algorithm. The aim of generating the SVD-based data and the HODMD-based data is to address the difficulty in the medicine field to elaborate a varied database with high quality enough to properly train deep learning algorithms, as implying high costs and very hard specialized work. Also note that the training, validation and test sets comprise different sequences, which are the ones that both achieve the splitting scheme proposed in this work and a reasonable balance in the number of samples in each class.
@@
Table 2: Summary of the combinations for $D_{train}$ used in the Cardiac Database Creation stage.
@@
Case Original Reconstructed Modes Reconstructed Modes Reconstructed

### v2

@@ 2.2 Heart Failure Prediction
In the second phase, Modal Decomposition-based Data Transform, the same process as in the Modal Decomposition-based Data Generation phase described in Subsection 2.1 is performed. This is, images with less noise and more discriminative features than the original echocardiography data are obtained by sequentially applying the SVD and the HODMD algorithms in the same way as previously described, increasing the differentiation between the heart states and improving the heart failure prediction performance. Therefore, this phase only acts as a feature extractor. However, unlike the Modal Decomposition-based Data Generation phase described in Subsection 2.1, the resulting images from this phase, Modal Decomposition-based Data Transform, are taken as input of the deep neural network for the prediction of the time of a heart failure of the corresponding sequence. On the other hand, the Modal Decomposition-based Data Generation phase leverages the output of each applied algorithm to enlarge the training database and form $\boldsymbol{D}_{train}$, directly fed to the deep neural network for training, containing more discriminative samples for the differentiation between the heart states, and addressing the limited availability of high-quality samples usual in the medicine field. Different types of data resulting from the Modal Decomposition-based Data Transform phase have been tested as input of the deep neural network for the prediction of heart failure times in Section 4: the original echocardiography images (that is, without the Modal Decomposition-based Data Transform), the reconstructions, or modes obtained after applying the SVD or the HODMD algorithm. In particular, according to the results (see Section 4), adopting HODMD data from the test video sequences of echocardiography images as the output of the Modal Decomposition-based Data Transform phase tends to give better heart failure time prediction performances. This demonstrates the efficacy of these techniques to extract more discriminative features associated with the different heart states from echocardiography images and to filter out the high noise inherent, leading to more accurate heart failure time predictions.
@@
The third phase, Deep Neural Network-based Heart Failure Prediction, computes the heart failure time prediction for each image generated in the previous phase, Modal Decomposition-based Data Transform. For this purpose, a deep neural network architecture based on the ViT is used. It has been trained using a self-supervised learning approach which improves the heart failure time prediction performance, even with scarce datasets (Das et al. (2024)). To the best of the authors’ knowledge, self-supervised learning has been barely explored in the related literature about heart failure prediction in echocardiography images, even less Masked Autoencoders (MAE) (He et al. (2022)). Moreover, standard architectures have been proposed as the general trend with no specific adaptations nor approaches to deal with this typical scenario in the medicine field of having a low number of high-quality samples. The training approach used involves two tasks: the Self-supervised Auxiliary Task, which aids in training the ViT for the regression task about heart failure time prediction, precisely the one of interest. This allows to increase the locality inductive bias of the ViT, and so to improve the performance on heart failure prediction in the usual scenario of scarce datasets. In addition, this training approach further reduces the dependence on very large training databases with high quality, which are costly to elaborate and which require much hard specialized work. As will be seen in Subsection 4.2, the proposed deep neural network trained with this self-supervised learning approach achieves a better performance than several ViT architectures and Convolutional Neural Networks (CNN), allowing to reduce computational resources.
@@
The architecture is represented in Fig. 5. Two tasks are jointly learned at the same time: the Self-supervised Auxiliary Task (Fig. 5 (b)), and the Regression Task (Fig. 5 (c)). The first one is aimed to reconstruct the missing patches from masked images, improving the discriminating capability of the deep neural network aimed to predict heart failure times, i.e., the one devoted to the Regression Task. Precisely, the weights of the Transformer Encoders used for both tasks are shared (highlighted in red in Fig. 5), allowing this joint training. Note that any self-supervised learning approach can be used, but the Masked Autoencoder (MAE) (He et al. (2022)) has been employed due to its popularity and superior performance (Das et al. (2024)). This training approach based on joint learning is different to the usual approach adopted in MAEs (He et al. (2022)), which first performs the pretraining of the ViT for reconstruction of the masked patches, and then the fine-tuning, using the same dataset, for the task of interest. This different training approach not only improves the discriminating capability and feature learning of the deep neural network, but also allows to speed up the training process by learning both tasks at the same time. The common input for the two tasks is an image representing a heart in a determined state (Fig. 5 (a)). As a result of the previous phase, Modal Decomposition-based Data Transform, this input image can be a reconstruction, or a mode obtained with the SVD and the HODMD algorithms. This fact gives the deep neural network the advantage of being able to directly use images as input, in the form of original echocardiography data, reconstructions, or modes obtained with the modal decomposition techniques. However, as said previously, using data obtained with the HODMD algorithm as input for the deep neural network (and so as the output of the Modal Decomposition-based Data Transform) has led to the best performance. The output is a prediction which represents the estimated time in which a heart failure will happen (Fig. 5 (d)). The ViT, as a deep neural network, is the one which addresses the non-linear relationships which could be present between the input and the output, in this case, between echocardiography images and heart failure times, regardless of whether the SVD and the HODMD algorithms have been employed or not. However, as will be seen in Section 4, the use of the SVD and HODMD algorithms for both feature extraction and data augmentation, and the fact that the proposed ViT is specifically adapted to scarce databases of echocardiography images, allow this proposed deep neural network to effectively model the relationship existing between echocardiography images and heart failure times. In addition, as already seen in Subsubsection 2.1.1, the HODMD algorithm, formulated with linear operators, is an approximation of the Koopman operator, allowing to deal with the non-linearities existing in the echocardiography imaging.
@@
The processing of the input (Fig. 5 (a)) for the Regression Task (the one of interest, Fig. 5 (c)), for heart failure time prediction (Fig. 5 (d)) closely follows the standard ViT architecture. That is, a division in non-overlapping patches is firstly performed. Next, the spatial dimension of the patches is flattened, followed by a normalization layer and a linear projection. Before processing the patches with the Transformer Encoder, positional information and a regression token, with a similar role to the class token, are added.
@@
Figure 5: Architecture of the proposed deep neural network with the training approach based on self-supervised learning. (a) The input: an image from an echocardiography video sequence, in the form of an original echocardiography sample, a mode, or a reconstruction obtained with the SVD and the HODMD algorithms, result from the Modal Decomposition-based Data Transform phase. (b) The Self-supervised Auxiliary Task, aimed to reconstruct the missing patches from the masked image, aiding in training the ViT for the regression task. (c) The Regression Task, the one of interest, for heart failure time prediction. (d) The predicted heart failure time of the input image.
@@
The backbone which processes the visual tokens is based on a Transformer Encoder, whose weights are shared with the Transformer Encoder used for the Self-supervised Auxiliary Task, learned at the same time (highlighted in red in Fig. 5). This has been done following the configuration of the ViT-T (Tiny) architecture, the shallowest one, to minimize overfitting (Das et al. (2024)), which is as follows: twelve stacked Transformer blocks (so $B_{E}=12$). Each Transformer block includes a multi-head attention layer, with three heads, projection dimension $192$, and without dropout.
@@
An empty branch (skipped connection) and dense layers with two units follow the multi-head attention layer, each one composed of a fully connected layer with the Gaussian Error Linear Unit (GELU) activation function. The use of skip connections addresses the vanishing gradient problem, inherent to deep neural networks, to considerably improve training convergence. The ratio of the dense layers used is $4$, that is, the number of hidden features is four times the amount of input features. The Transformer block finally introduces a skip connection, connecting the output of the previous one to the output of these dense layers. After all the $B_{E}$ Transformer blocks, features are normalized and introduced into a regression head, whose output is the estimation of the time in which a heart failure will happen (Fig. 5 (d)), according to the input image (Fig. 5 (a)).
@@
Regarding the Self-supervised Auxiliary Task (Fig. 5 (b)), the processing is based on the MAE (He et al. (2022)). That is, first, patches from the input image (Fig. 5 (a)) are randomly masked before the tokenization. After that, the same Transformer Encoder as the one for the Regression Task is used. Precisely, the weights between both Transformer Encoders are shared (highlighted in red in Fig. 5). Therefore, during training, these weights are adapted to both tasks, improving the discriminating capability of the ViT used for heart failure time prediction, even with scarce databases. After the $B_{E}$ Transformer blocks, a Transformer Decoder is used for the reconstruction of the missing patches, much shallower than the Encoder. This Decoder has been designed by first incorporating a decoder embedding which adapts the projection dimension to $128$ via a fully connected layer. After adding the masked tokens and the position embedding, two stacked Transformer blocks have been used ($B_{D}=2$). Each of these blocks includes a multi-head attention layer with $16$ heads. Similarly, empty branches and dense layers with two units and a ratio of 4, each one with the GELU activation function, have also been employed. After all the $B_{D}$ Transformer blocks, features are normalized, the regression token is removed, and a decoder head is used for the reconstruction of the missing patches.
@@
Considering the joint learning on both tasks at the same time, the final convex loss function is as follows:
@@
where $L_{reg}$ is the regression loss between the estimated heart failure time and the real one; $L_{SSAT}$ is the loss component between the original and reconstructed image, computed only for the masked patches (He et al. (2022)), and $\alpha$ is the loss scaling factor, weighing the importance of each task. Both loss functions are based on the Mean Square Error (MSE). Note that, after training, like in the usual approach of MAEs (He et al. (2022)), for the Deep Neural Network-based Heart Failure Prediction phase, only the branch concerning the Regression Task (Fig. 5 (c)) is used, and the branch about the Self-supervised Auxiliary Task (Fig. 5 (d)) is discarded.
@@
The deep neural network uses the annotated cardiac database of images, $\boldsymbol{D}_{train}$, obtained in the Cardiac Database Creation stage described in Subsection 2.1 for training. Concerning the optimization technique, the AdamW algorithm has been used. The employed learning rate strategy is based on the warm-up cosine policy (Lee et al. (2021)) that computes the learning rate at iteration $i$ by the expression $\lambda_{i}=0.5\times\lambda_{t}\times(1+cos(\pi\times(i-N_{w})/(N_{iter}-N_{w})))$, where $N_{iter}$ is the maximum number of iterations. Following the configurations used in Das et al. (2024), the target learning rate is fixed to $\lambda_{t}=2.5\mathrm{e}{-4}$, and the warmup steps $N_{w}$ to $5$. The momentum and weight decay are set to 0.9 and 0.05, respectively. The batch size has been set to 64 images due to the limitations of the physical memory of the GPUs available (Nvidia A100 and Nvidia RTX A4500). About the database division, a splitting scheme of $60\%-20\%$ for training and validation, respectively, has been adopted, balancing the cardiac categories to ensure that every set contains approximately the same number of samples for each heart state. The batch generation scheme includes data augmentation techniques based on geometric transformations (namely: resizing, random horizontal flips, and random erasing), applied to the echocardiography data. In this way, the model becomes more robust to different perspectives acquired of the heart. In addition, these data augmentation techniques complement to the Modal Decomposition-based Data Generation phase already explained in Subsection 2.1, marginally improving the performance compared to the modal decomposition.
@@
In the fourth and last phase, Fusion of Heart Failure Predictions, an estimation of the time of happening a heart failure in the input test sequence is determined by combining multiple predictions. This process is illustrated in Fig. 6. Each image $i$ from the $N+1$ images depicted were obtained from the input test sequence according to the process explained in the second phase, Modal Decomposition-based Data Transform. That is, these images can be either original echocardiography images, or modes or reconstructions obtained with the SVD and HODMD modes. Then, each of these images is processed by the deep neural network as explained in the third phase, Deep Neural Network-based Heart Failure Prediction. Therefore, each of these images has associated a time prediction $P_{i}$, that is, the estimation of the age of the patient in which a heart failure will happen. The set of time predictions $\{P_{i}\}$ is then used to compute a unique time prediction associated to the input test sequence $\hat{P}$. For this purpose, the average has been used. As a result, an estimation of the time in which a heart failure will happen is obtained. Note that relying on a sequence of echocardiography images instead of a single image to predict the time of heart failures is more robust, as using more data can help doctors to make more confident heart prognoses. In addition, certain heart pathologies might lead to arrhythmias or patterns in the heart cycles which, therefore, constitute temporal information which can be obtained from sequences, and which could be valuable to accurately predict heart failures.
@@ 3 Database
The heart failure prediction system here proposed has been tested and validated using a database of echocardiography images. It is composed of video sequences representing the following heart states: Control (CTL), Obesity (OB), and Systemic Hypertension (SH). This database has been elaborated in a collaboration with the Centro Nacional de Investigaciones Cardiovasculares (CNIC). A sample image from each of these heart states is shown in Fig. 7. As can be seen, echocardiography imaging inherently has much noise, as expected in this modality of data. In addition, different perspectives are taken, increasing the complexity of the heart failure prediction task. Only the area of the heart, i.e., the region of interest (ROI), is taken for the prediction system, as specified in the Data Homogenization phase described in Subsection 2.1. The obtained heart areas have different resolutions, showing the heterogeneity of the echocardiography imaging: in mean and standard deviation, the spatial resolution is $(675.68\pm 47.28)\times(583.15\pm 0.49)$ pixels. However, the aspect factor among the images barely changes and is close to 1, so these are practically square: $1.16\pm 0.08$. Therefore, resizing to square images for the deep neural networks barely deforms the heart areas, and does not suppose a degradation of the heart failure prediction performance, as preserving the spatial features characterizing the different heart states.
@@
A summary of the main characteristics of the cardiac database is presented in Table 1. It includes information about the times in which heart failures happen. An average and a standard deviation of the age of the heart failure event are represented for each heart state. As can be seen, unhealthy classes (OB, and SH) tend to have earlier heart failures than in the case of healthy hearts (CTL), as expected. This cardiac database is composed of the original echocardiography images, and the associated SVD-based data and the HODMD-based data. These data have been generated according to the process made in the Modal Decomposition-based Data Generation phase described in Subsection 2.1. That is, the SVD-based data were obtained by applying the SVD algorithm on each video sequence of echocardiography images. For each of these sequences, reconstructions and modes were obtained with the SVD algorithm (denoted as $\boldsymbol{D}_{SVD1\_R}$ and $\boldsymbol{D}_{SVD1\_M}$, respectively, following the notation from Subsection 2.1). After that, the HODMD-based data were generated by applying the HODMD algorithm on each sequence, but using their associated SVD reconstructions instead of the original images. As a result, for each sequence, reconstructions and modes were obtained ($\boldsymbol{D}_{HODMD\_R}$, $\boldsymbol{D}_{HODMD\_M}$, respectively). Finally, the SVD algorithm was applied on each sequence, using their associated HODMD modes, taking in this case the reconstructions ($\boldsymbol{D}_{SVD2\_R}$). In Table 1, the column ’# Snapshots’ indicates the total number of original snapshots of the sequences. Note that this number equals the number of associated reconstructed ones after applying the SVD algorithm, and also the number of reconstructions obtained after using the HODMD algorithm. The columns ’# SVD modes’ and ’# HODMD modes’ represent the total number of SVD modes and of HODMD modes obtained, respectively. The aim of generating the SVD-based data and the HODMD-based data is to address the difficulty in the medicine field to elaborate a varied database with high quality enough to properly train deep learning algorithms, as implying high costs and very hard specialized work. Also note that the training, validation and test sets comprise different sequences, which are the ones that both achieve the splitting scheme proposed in this work and a reasonable balance in the number of samples in each class.
@@
Table 2: Summary of the combinations for $\boldsymbol{D}_{train}$ used in the Cardiac Database Creation stage.
@@
Case $\boldsymbol{D}_{orig}$ $\boldsymbol{D}_{SVD1\_R}$ $\boldsymbol{D}_{SVD1\_M}$ $\boldsymbol{D}_{HODMD\_R}$ $\boldsymbol{D}_{HODMD\_M}$ $\boldsymbol{D}_{SVD2\_R}$